OpenStudy (anonymous):

z varies inversely with the product of x and y if z=36 when x=4 and y=2

OpenStudy (whpalmer4):

Do you know what inverse variation means?

OpenStudy (anonymous):

i need to figure out how to set it up then plug in the numbers and i cant find out how to set it up

OpenStudy (whpalmer4):

Yes, but the first thing is to understand the relationship: when someone says "x varies inversely with y", what does that mean?

OpenStudy (whpalmer4):

choose one of these two options: 1) y = kx 2) y = k/x (or xy = k)

OpenStudy (anonymous):

i tried to use z=kx/y and that was not the right one

OpenStudy (whpalmer4):

you need to use \[z = \frac{k}{xy}\]because \(z\) varies inversely with the product of \(xy\)